OpenStudy (anonymous):

solve the equation. 6(3x-2)=15+(x+7)

OpenStudy (anonymous):

18x-12=15+(x+7) 18x-12-15=x+7 18x-27-7=x 18x-34=x -34=(1)x-18x -34=-17x 34=17x 34/17=x x=2

OpenStudy (anonymous):

step by step for ya lol